# International school for advanced studies

### v1

@@ Acknowledgements.
Eugenio Montale, Ossi di Seppia (1925).
@@
Immanuel Kant, Critique of Pure Reason (1781)
@@ 1 Introduction
Learning complex functions defined in high-dimensional spaces is an extremely hard task because the number of samples must grow exponentially with the dimension of the ambient space [73, 74]. Luckily most real-world data sets are characterized by regularities that induce strong constraints and correlations between the features of the data. In natural images, nearby pixels have strongly correlated colors and intensities and often have few degrees of freedom [75]. Regularities occurring in text and in the sound of spoken letters [76, 77] similarly result in a low number of relevant variations. In molecular dynamics simulations, the region of the configuration space visited by a complex molecule of $N$ atoms has a dimensionality much smaller than the $3N$ degrees of freedom available, due to the constraints imposed by the bonds between the atoms and the steric hindrance between parts of the molecule [65]. The number of possible variations in proteins belonging to the same family is much smaller than the sequence length [78] because of the conserved amino acids at some positions of their sequences and correlated variation in others. Therefore, despite real-world data possessing a very large number of features, in practice, they often tend to concentrate around low-dimensional, often highly non-linear, manifolds. The dimension of the manifold is referred to as intrinsic dimension (ID) of the data set. In many circumstances the knowledge of the ID also gives valuable information about the inner properties of the dataset [79, 80, 81].
@@ 2 The Gride estimator
Maximizing Eq. (12) we get an intrinsic dimension estimator for $d$ that generalizes the TwoNN estimator to higher order ratios. For this reason, we call our ID estimator Generalized Ratios Intrinsic Dimension Estimator, in short, Gride.
@@ 1 Simulated data sets
We first describe the scale analysis of the ID with Gride on $N_{tot}=32000$ data points sampled from a uniform distribution on $d$-hypercubes where $d=\{2,5,10,20,\\ 50\}$ (Fig. 3. To increase the scale, we double each time the range $(k_{1},k_{2})$ of nearest neighbors’ ranks: $(k_{1},k_{2})=\{(1,2),(2,4),(4,8),$...\} and plot the estimated ID as a function of $N_{tot}/\overline{k}$ with $\overline{k}=\nicefrac{{(k_{1}+k_{2})}}{{2}}$. $N_{tot}/\overline{k}$ plays the role of the "scale parameter": the higher $N_{tot}/\overline{k}$ the smaller on average the size of the neighborhood.
@@
Figure 4 shows the scale analysis of the ID on $N_{tot}=16000$ points sampled from a 1-$d$ spiral, a 2-$d$ standard normal distribution and 2-$d$ Moebius strip. The embedding dimensions $D$ are 3 (a-b) and 6 (c). Gaussian noise of standard deviation $\sigma=\nicefrac{{0.01}}{{\sqrt{D}}}$ is independently added to each coordinate of the manifolds. At large $N/\overline{k}$, the typical distance between a point and its $k$-nearest neighbors is smaller then the noise standard deviation $\sigma$, and all the methods overestimate the ID of the underlying manifold. Increasing the neighborhood size, when the typical distance becomes larger than $\sigma$ the estimated IDs approach the true dimension of the manifolds. However, the rate and the scale at which the ID reaches the ground truth value are different for each estimator, and, in all cases, Gride stabilizes around $d$ at smaller scales than the competitors.
More precisely, for the 1dd spiral (a), the noise standard deviation $\sigma\sim 6.0\cdot 10^{-3}$ and is similar to the average distance to the first two nearest neighbors and $N/\overline{k}\sim 10^{4}$ corresponds to distances of the order of $3\sigma$. TwoNN and Gride are the only estimators able to evaluate the ID at very local scales $N/\overline{k}>10^{4}$ where the ID is almost 3, equal to the embedding dimension. DANCo, ESS, and MLE reach $N/\overline{k}\sim 10^{4}$ where the ID is close to 3. The GeoMLE profile instead starts from $N/\overline{k}\sim 2\cdot 10^{3}$ a scale at which the ID is around 1.9. In the range $10^{2}<N/\overline{k}<10^{3}$, the Gride profile decreases faster than those of the other estimators and reaches 1.006 around $N/\overline{k}\sim 5\cdot 10^{2}$. At the same scale, the IDs measured by TwoNN, DANCo, ESS, MLE, and GeoMLE are $[1.3,1.32,1.4,1.48,1.35]$ respectively. Increasing further the neighborhood size ($N/\overline{k}<5\cdot 10^{2}$), Gride remains constant and equal to one and the other estimators gradually approach the same value with the exception of ESS which shows a small increase for $N/\overline{k}\sim 10^{1}$. In this range, $\hat{d}$ equals the true ID of the manifold.
@@
TwoNN in addition to reaching the ground truth ID at larger scales than Gride, in the case of the normal distribution (b) does not show a plateau at all; DANCo measures very stable ID profiles over a wide range of scales; ESS tends to overestimate the IDs with plateaus that are on average shorter than those of Gride, TwoNN, DANCo and MLE;
Figure 6: Simulated data sets without additional noise. The figure shows the ID as a function of different neighborhood sizes on ten simulated data sets without noise. The manifold where the points lie is written on the top of each panel and the analytic equations are reported in Fig. 4 and Fig. 5. We compare six estimators as reported in the legend of the first panel. We plot the neighborhood size on the x-axis in logarithmic scale as a function of $\nicefrac{{N}}{{\overline{k}}}$: NN is the data set size and $\overline{k}$ is the neighborhood range.
MLE is very similar to TwoNN and Gride. It is more robust than TwoNN because it considers a larger set of points to infer the ID (e.g $k=10$, see [96]) but its estimates are less local and larger than those of TwoNN and Gride for a given scale; GeoMLE is most unstable likely due to the large neighborhood size used to fit the ID. shows the analysis of six other types of manifolds. For comparison, Fig. 6 shows the same results without additional noise.
@@ 2 Real data sets
On MNIST digit one (Fig. 7-a), the ID increases with $N/\overline{k}$ from 8 to 13, for all the estimators. Gride, TwoNN, MLE, and GeoMLE are consistent with each other in the range $100<N/\overline{k}<700$, while ESS and DANCo estimate a slightly larger ID. This example shows the importance to compare the IDs on the same "scale" given by $N/\overline{k}$. For instance, at $N/\overline{k}\sim 700$, Gride, TwoNN, MLE and GeoMLE measure the ID on data sets sizes equal to 6742, 1685, 3371, and $6742\cdot 0.632\sim 4260$ respectively, but the estimated ID is similar for all of them.
@@ 3 Computational cost of Gride vs other ID estimators
To compare the speed of the estimators we measure how the execution time grows as a function of the number of samples $N$ and of the number of input coordinates $P$ using the CIFAR10 data set [108] which consists of 50000 color images of size $3\times 32\times 32$. We vary $N$ taking random subsets of size $\{N,N/2,N/4,...N/64\}$, and vary $P$ by resizing the resolution of 5000 images (those of the cats) from $3\times 16\times 16$ ($P=768$) to $3\times 181\times 181$ ($P=98283)$. Each estimator is evaluated once on each data size. The hyperparameters are the same as those used in Sec. 1 and 2: $k\leq 10$ for DANCo, ESS and MLE, $5\leq k\leq 15$ for GeoMLE. For TwoNN by definition $k\leq 2$, and for Gride we consider $k\leq 64$. The computation is done on an Intel Xeon E5-2690v3 processor with 12 cores and 24 threads.
@@
Figure 8: Time benchmarks for different estimators We compare the computational time of the estimators reported in the legend as a function of the data set size NN a:, and the number of features $P$ b: of each data point. The data set is CIFAR10. We vary NN choosing a random subset of increasing size, keeping $P=1024$. We vary $P$ increasing the resolution of $N=5000$ images.
In both cases, Gride, TwoNN and MLE are always one order of magnitude faster than DANCo and GeoMLE, and almost two orders of magnitude faster than ESS. When $N=50000$ (a) the ID evaluation takes 2 minutes for Gride, TwoNN and MLE, 19 and 26 minutes for DANCo and GeoMLE (x10/15 speedup), 1 hour and 20 minutes for ESS (x40 speedup). The comparison is even more favorable when $P$ is large (Fig. 8-b): when $P\sim 10^{4}$ the execution time is 3.3 seconds for Gride TwoNN and MLE, 60 and 80 seconds for DANCo and GeoMLE (x20/25 speedup), almost 10 minutes for ESS (x160 speedup).
@@
The high computational cost is also the reason why we chose not to increase the neighborhood size by adjusting the nearest neighbor order $k$ in DANCo, ESS. In ESS, increasing $k$ from 10 to 30, the computational time becomes three times larger than what is shown in Fig. 8. In MLE (and GeoMLE), from a computational time perspective, it is feasible to increase $k$ to evaluate the ID on multiple scales. However, as we saw in the previous section, GeoMLE is very unstable for high $k$ values especially when the number of data is small. MLE has a memory overhead with respect to Gride since all the neighbors up to $k$ are needed to compute the ID (see Eq. (3, 4)) and the memory scales as $N\cdot k$. In Gride the memory grows as $N\cdot\log_{2}k$, because only the neighbors’ ranks of an integer power of two are stored.
@@ 6 Discussion
In this chapter, we described Gride, an extension of the TwoNN estimator to efficiently estimate the ID on multiple scales. Gride is orders of magnitude faster than other nearest neighbor-based estimators (DANCo, GeoMLE, ESS), does not depend on any hyperparameter because it evaluates the IDs on multiple scales and does not require to decimate the data set like TwoNN. On data sets with noise, Gride allows exiting the "noisy region" at shorter length scales making the ID easier to identify.
@@
The estimator is efficient and reliable, but some issues related to its properties are still open. The first issue is related to the error estimated for large nearest-neighbor ranks. The variance of the Gride likelihood as a function of $d$ is a decreasing function of $k$. As a result, increasing $k$ the observed information grows and the estimated confidence intervals become small (see Eq. (15)). Figure 9-a shows the width of 95% confidence intervals estimated on the 2-dimensional uniform distribution analyzed in Fig. 4-a as a function of $k$.
@@
Figure 9: Confidence intervals of the Gride estimator. In panel a we plot the 95% large-sample confidence intervals (CI) of the Gride estimator for an increasing value of $k_{1}$. As $k_{1}$ grows the (CI) becomes small as a result of the decrease of the variance (see Fig. 2). Panel b shows the cumulative distribution functions (CDF) of the Gride likelihood.
@@
When $k=256$ the estimated ID with its confidence is $1.930\pm 0.001$. However, when the size of the neighborhood is large, the Poisson process assumption used to derive Gride is violated more often, mainly due to the finite size of the data set. This is an unfortunate case where the model becomes more confident when the prediction is wrong.
@@
To overcome this difficulty one should always perform an additional model assessment step. A visual inspection of the deviation of the empirical cumulative distribution (CDF) of the Gride likelihood (Eq. (12)) from the analytical one, gives a first, qualitative, hint about the consistency between the model and the data. In Fig. 9-b we compare the empirical CDF (blue) with the analytical one based on the ground truth (ID = 2, orange) and the estimated ID for $k=64$ and $k=256$. As expected, the deviation from the ground truth is larger for higher values of $k$ and, in this setting, concerns the tail where the $\mu_{k}$ are larger which is more populated in the empirical CDF. In general, the higher the ID, the smaller the $\mu_{k}=\nicefrac{{r_{2k}}}{{r_{k}}}$ because in high dimension is more likely for nearest neighbors to occupy regions of space at a similar distance from a given point (see also Fig. 2). The presence of boundaries limits portions of space that could be occupied by data points, making the higher tail of the likelihood fatter and the estimated ID smaller. Unfortunately, it is hard to make quantitative statements about the goodness of the fit in the general case but the inspection can nonetheless rule out cases where the fit is manifestly wrong.
@@
A second limitation of Gride is the assumption of constant ID for all the data points. This is a common hypothesis for most ID estimators [64], but [109] showed that in real-world data sets the ID is in general not constant, and being able to identify the homogeneous ID components can add valuable information about the data. To segment the data according to the local ID, [109] proposed Hidalgo, a Bayesian mixture model based on the TwoNN Pareto likelihood Eq. (1). A difficulty of Hidalgo is related to the large overlap of the Pareto likelihoods (see Fig 2-a) which makes a single $\mu_{i}$ compatible with models of different IDs. The Gride likelihoods at relatively low neighbor ranks (eg $k_{1}=4$, $k_{2}=8$) are less overlapped and could be used in place of the Pareto ones to improve the discriminative power of Hidalgo. A promising future direction of research would be to combine Hidalgo and Gride for a multiscale ID estimation where the intrinsic dimension is not constant in the data set.

### v2

@@ Acknowledgements.
Eugenio Montale, Ossi di Seppia, 1925.
@@
Immanuel Kant, Critique of Pure Reason, 1781.
@@ 1 Introduction
Learning complex functions defined in high-dimensional spaces is an extremely hard task because the number of samples must grow exponentially with the dimension of the ambient space [73, 74]. Luckily most real-world data sets are characterized by regularities that induce strong constraints and correlations between the features of the data. In natural images, nearby pixels have strongly correlated colors and intensities and often have few degrees of freedom [75]. Regularities occurring in text and in the sound of spoken letters [76, 77] similarly result in a low number of relevant variations. In molecular dynamics simulations, the region of the configuration space visited by a complex molecule of $N$ atoms has a dimensionality much smaller than the $3N$ degrees of freedom available, due to the constraints imposed by the bonds between the atoms and the steric hindrance between parts of the molecule [65]. The number of possible variations in proteins belonging to the same family is much smaller than the sequence length [78] because of the conserved amino acids at some positions of their sequences and correlated variation in others. Therefore, despite real-world data possessing a very large number of features, in practice, they often tend to concentrate around low-dimensional, often highly non-linear, manifolds. The dimension of the manifold is referred to as intrinsic dimension (ID) of the data set. In many circumstances, the knowledge of the ID also gives valuable information about the inner properties of the dataset [79, 80, 81].
@@ 2 The Gride estimator
By maximizing Eq. (12), we get an intrinsic dimension estimator for $d$ that generalizes the TwoNN estimator to higher order ratios. For this reason, we call our ID estimator Generalized Ratios Intrinsic Dimension Estimator, in short, Gride.
@@ 1 Simulated data sets
We first describe the scale analysis of the ID with Gride on $N_{tot}=32000$ data points sampled from a uniform distribution on $d$-hypercubes where $d=\{2,5,10,20,\\ 50\}$ (Fig. 3. To increase the scale, we double each time the range $(k_{1},k_{2})$ of nearest neighbors’ ranks: $(k_{1},k_{2})=\{(1,2),(2,4),(4,8),$...\} and plot the estimated ID as a function of $N_{tot}/\overline{k}$ with $\overline{k}=\nicefrac{{(k_{1}+k_{2})}}{{2}}$. $N_{tot}/\overline{k}$ plays the role of the "scale parameter": the higher $N_{tot}/\overline{k}$, the smaller on average the size of the neighborhood.
@@
Figure 4 shows the scale analysis of the ID on $N_{tot}=16000$ points sampled from a 1-$d$ spiral, a 2-$d$ standard normal distribution, and a 2-$d$ Moebius strip. The embedding dimensions $D$ are 3 (a-b) and 6 (c). Gaussian noise of standard deviation $\sigma=\nicefrac{{0.01}}{{\sqrt{D}}}$ is independently added to each coordinate of the manifolds. At large $N/\overline{k}$, the typical distance between a point and its $k$-nearest neighbors is smaller then the noise standard deviation $\sigma$, and all the methods overestimate the ID of the underlying manifold. Increasing the neighborhood size, when the typical distance becomes larger than $\sigma$ the estimated IDs approach the true dimension of the manifolds. However, the rate and the scale at which the ID reaches the ground truth value are different for each estimator, and, in all cases, Gride stabilizes around $d$ at smaller scales than the competitors.
More precisely, for the 1dd spiral (a), the noise standard deviation $\sigma\sim 6.0\cdot 10^{-3}$ and is similar to the average distance to the first two nearest neighbors and $N/\overline{k}\sim 10^{4}$ corresponds to distances of the order of $3\sigma$. TwoNN and Gride are the only estimators able to evaluate the ID at very local scales $N/\overline{k}>10^{4}$ where the ID is almost 3, equal to the embedding dimension. DANCo, ESS, and MLE reach $N/\overline{k}\sim 10^{4}$ where the ID is close to 3. The GeoMLE profile instead starts from $N/\overline{k}\sim 2\cdot 10^{3}$, a scale at which the ID is around 1.9. In the range $10^{2}<N/\overline{k}<10^{3}$, the Gride profile decreases faster than those of the other estimators and reaches 1.006 around $N/\overline{k}\sim 5\cdot 10^{2}$. At the same scale, the IDs measured by TwoNN, DANCo, ESS, MLE, and GeoMLE are $[1.3,1.32,1.4,1.48,1.35]$ respectively. Increasing further the neighborhood size ($N/\overline{k}<5\cdot 10^{2}$), Gride remains constant and equal to one and the other estimators gradually approach the same value with the exception of ESS which shows a small increase for $N/\overline{k}\sim 10^{1}$. In this range, $\hat{d}$ equals the true ID of the manifold.
@@
TwoNN in addition to reaching the ground truth ID at larger scales than Gride, in the case of the normal distribution (b) does not show a plateau at all; DANCo measures very stable ID profiles over a wide range of scales; ESS tends to overestimate the IDs with plateaus that are on average shorter than those of Gride, TwoNN, DANCo, and MLE;
Figure 6: Simulated data sets without additional noise. The figure shows the ID as a function of different neighborhood sizes on ten simulated data sets without noise. The manifold where the points lie is written on the top of each panel, and the analytic equations are reported in Fig. 4 and Fig. 5. We compare six estimators as reported in the legend of the first panel. We plot the neighborhood size on the x-axis in logarithmic scale as a function of $\nicefrac{{N}}{{\overline{k}}}$: NN is the data set size and $\overline{k}$ is the neighborhood range.
MLE is very similar to TwoNN and Gride. It is more robust than TwoNN because it considers a larger set of points to infer the ID (e.g $k=10$, see [96]), but its estimates are less local and larger than those of TwoNN and Gride for a given scale; GeoMLE is most unstable likely due to the large neighborhood size used to fit the ID. shows the analysis of six other types of manifolds. For comparison, Fig. 6 shows the same results without additional noise.
@@ 2 Real data sets
On MNIST digit one (Fig. 7-a), the ID increases with $N/\overline{k}$ from 8 to 13, for all the estimators. Gride, TwoNN, MLE, and GeoMLE are consistent with each other in the range $100<N/\overline{k}<700$, while ESS and DANCo estimate a slightly larger ID. This example shows the importance to compare the IDs on the same "scale" given by $N/\overline{k}$. For instance, at $N/\overline{k}\sim 700$, Gride, TwoNN, MLE, and GeoMLE measure the ID on data sets sizes equal to 6742, 1685, 3371, and $6742\cdot 0.632\sim 4260$ respectively, but the estimated ID is similar for all of them.
@@ 3 Computational cost of Gride vs other ID estimators
To compare the speed of the estimators, we measure how the execution time grows as a function of the number of samples $N$ and of the number of input coordinates $P$ using the CIFAR10 data set [108] which consists of 50000 color images of size $3\times 32\times 32$. We vary $N$ taking random subsets of size $\{N,N/2,N/4,...N/64\}$, and vary $P$ by resizing the resolution of 5000 images (those of the cats) from $3\times 16\times 16$ ($P=768$) to $3\times 181\times 181$ ($P=98283)$. Each estimator is evaluated once on each data size. The hyperparameters are the same as those used in Sec. 1 and 2: $k\leq 10$ for DANCo, ESS and MLE, $5\leq k\leq 15$ for GeoMLE. For TwoNN by definition $k\leq 2$, and for Gride we consider $k\leq 64$. The computation is done on an Intel Xeon E5-2690v3 processor with 12 cores and 24 threads.
@@
Figure 8: Time benchmarks for different estimators We compare the computational time of the estimators reported in the legend as a function of the data set size NN a:, and the number of features $P$ b: of each data point. The data set is CIFAR10. We vary NN choosing a random subset of increasing size, keeping $P=1024$. We vary $P$ by increasing the resolution of $N=5000$ images.
In both cases, Gride, TwoNN, and MLE are always one order of magnitude faster than DANCo and GeoMLE, and almost two orders of magnitude faster than ESS. When $N=50000$ (a) the ID evaluation takes 2 minutes for Gride, TwoNN and MLE, 19 and 26 minutes for DANCo and GeoMLE (x10/15 speedup), 1 hour and 20 minutes for ESS (x40 speedup). The comparison is even more favorable when $P$ is large (Fig. 8-b): when $P\sim 10^{4}$, the execution time is 3.3 seconds for Gride TwoNN and MLE, 60 and 80 seconds for DANCo and GeoMLE (x20/25 speedup), almost 10 minutes for ESS (x160 speedup).
@@
The high computational cost is also the reason why we chose not to increase the neighborhood size by adjusting the nearest neighbor order $k$ in DANCo, ESS. In ESS, increasing $k$ from 10 to 30, the computational time becomes three times larger than what is shown in Fig. 8. In MLE (and GeoMLE), from a computational time perspective, it is feasible to increase $k$ to evaluate the ID on multiple scales. However, as we saw in the previous section, GeoMLE is very unstable for high $k$ values, especially when the number of data is small. MLE has a memory overhead with respect to Gride since all the neighbors up to $k$ are needed to compute the ID (see Eq. (3, 4)) and the memory scales as $N\cdot k$. In Gride the memory grows as $N\cdot\log_{2}k$, because only the neighbors’ ranks of an integer power of two are stored.
@@ 6 Discussion
In this chapter, we described Gride, an extension of the TwoNN estimator to efficiently estimate the ID on multiple scales. Gride is orders of magnitude faster than other nearest neighbor-based estimators (DANCo, GeoMLE, ESS), does not depend on any hyperparameter because it evaluates the IDs on multiple scales, and does not require to decimate the data set like TwoNN. On data sets with noise, Gride allows exiting the "noisy region" at shorter length scales, making the ID easier to identify.
@@
The estimator is efficient and reliable, but some issues related to its properties are still open. The first issue is related to the error estimated for large nearest-neighbor ranks. The variance of the Gride likelihood as a function of $d$ is a decreasing function of $k$. As a result, by increasing $k$, the observed information grows and the estimated confidence intervals become small (see Eq. (15)). Figure 9-a shows the width of 95% confidence intervals estimated on the 2-dimensional uniform distribution analyzed in Fig. 4-a as a function of $k$.
@@
Figure 9: Confidence intervals of the Gride estimator. In panel a we plot the 95% large-sample confidence intervals (CI) of the Gride estimator for an increasing value of $k_{1}$. As $k_{1}$ grows, the (CI) becomes small as a result of the decrease of the variance (see Fig. 2). Panel b shows the cumulative distribution functions (CDF) of the Gride likelihood.
@@
When $k=256$, the estimated ID with its confidence is $1.930\pm 0.001$. However, when the size of the neighborhood is large, the Poisson process assumption used to derive Gride is violated more often, mainly due to the finite size of the data set. This is an unfortunate case where the model becomes more confident when the prediction is wrong.
@@
To overcome this difficulty one should always perform an additional model assessment step. A visual inspection of the deviation of the empirical cumulative distribution (CDF) of the Gride likelihood (Eq. (12)) from the analytical one, gives a first, qualitative hint about the consistency between the model and the data. In Fig. 9-b, we compare the empirical CDF (blue) with the analytical one based on the ground truth (ID = 2, orange) and the estimated ID for $k=64$ and $k=256$. As expected, the deviation from the ground truth is larger for higher values of $k$ and, in this setting, concerns the tail where the $\mu_{k}$ are larger which is more populated in the empirical CDF. In general, the higher the ID, the smaller the $\mu_{k}=\nicefrac{{r_{2k}}}{{r_{k}}}$ because in high dimension it is more likely for nearest neighbors to occupy regions of space at a similar distance from a given point (see also Fig. 2). The presence of boundaries limits portions of space that could be occupied by data points, making the higher tail of the likelihood fatter and the estimated ID smaller. Unfortunately, it is hard to make quantitative statements about the goodness of the fit in the general case, but the inspection can nonetheless rule out cases where the fit is manifestly wrong.
@@
A second limitation of Gride is the assumption of constant ID for all the data points. This is a common hypothesis for most ID estimators [64], but [109] showed that in real-world data sets the ID is in general not constant, and being able to identify the homogeneous ID components can add valuable information about the data. To segment the data according to the local ID, [109] proposed Hidalgo, a Bayesian mixture model based on the TwoNN Pareto likelihood Eq. (1). A difficulty of Hidalgo is related to the large overlap of the Pareto likelihoods (see Fig. 2-a), which makes a single $\mu_{i}$ compatible with models of different IDs. The Gride likelihoods at relatively low neighbor ranks (e.g. $k_{1}=4$, $k_{2}=8$) are less overlapped and could be used in place of the Pareto ones to improve the discriminative power of Hidalgo. A promising future direction of research would be to combine Hidalgo and Gride for a multiscale ID estimation where the intrinsic dimension is not constant in the data set.